# Understanding Self-Supervised Learning via Gaussian Mixture Models

### v2

@@ 6 Multi-Modal Gaussian Mixtures and Self-Supervised Learning
In the previous section, we considered the setup in which a point and its augmentation follow the same distribution. This is a sound assumption for methods like SimCLR and SimSiam where augmentations of points (i.e. images) are defined by transformations (e.g., cropping, color jittering etc) on the image. This assumption may not hold in general. Text embeddings models like DPR Karpukhin et al. (2020) or image-text embedding models like CLIP Radford et al. (2021) consider input samples as a pair of points following (possibly) different distribution. For instance, each sample could be a pair of an image with its corresponding caption (for CLIP) or a search query with a relevant document that answers the query (for DPR). The objective of methods considering multi-modal data is to learn a joint representation space for both the modalities such that representations of points in a pair are close, while being far from random draws. These joint representations can be used downstream for finding similarities between points (belonging to different modalities). Surprisingly, representations learned for each modality are competitive to fully supervised representations Radford et al. (2021). Our goal is to theoretically analyze the representations learned by multi-modal models, specifically the CLIP model.
CLIP Gaussian Mixtures (CLIPGMM)
We define CLIPGMM to capture multi-modal data in a theory-friendly setting.
@@
(CLIPGMM) A CLIP Gaussian mixture (CLIPGMM) is defined as the probability distribution $F_{\mathrm{clip}}=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{V,k},{\bm{\Sigma}}_{V})\times\mathcal{N}({\bm{\mu}}_{T,k},{\bm{\Sigma}}_{T})$ where $w_{k}$ are the mixture weights, $\{{\bm{\mu}}_{V,k}\}_{k\in[K]},{\bm{\Sigma}}_{V}$ and $\{{\bm{\mu}}_{T,k}\}_{k\in[K]},{\bm{\Sigma}}_{T}$ are the parameters for the two coordinate spaces respectively.
CLIPGMM is hence a mixture of product distribution over Gaussians. To elaborate, a sampling $({{\bm{x}}}_{V},{{\bm{x}}}_{T})\sim F_{\mathrm{clip}}$ is a two step process. We first sample an underlying component index, and then draw independent samples from the component with the same index in the respective coordinate space. $F_{\mathrm{clip}}$ can be used to define marginal distribution over each coordinate space; $F_{V}=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{V,k},{\bm{\Sigma}}_{V})$. Marginal over $T$ space can be defined similarly.
Problem Setup
Following CLIP Radford et al. (2021), we want to learn two different representation functions (one for each modality). We let these functions be linear mappings (say, ${{\bm{A}}}_{V}\in\mathbb{R}^{d_{1}\times r},{{\bm{A}}}_{T}\in\mathbb{R}^{d_{2}\times r}$). Recall that optimal mapping matrices for each coordinate space would have their column space equal to the Fisher subspace for the marginal distributions $F_{V},F_{T}$. We show CLIP learns a mapping onto the subset of the Fisher subspace by jointly optimizing over multi-modal pairs, without augmentations.
@@ 6.1 Optimization Objective
We now formalize the CLIP InfoNCE loss Radford et al. (2021). Concretely, let the representation of a sample $({{\bm{x}}}_{V},{{\bm{x}}}_{T})$ be given by $({{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V},{{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})$, where ${{\bm{A}}}_{V},{{\bm{A}}}_{T}$ are the mappings for the two coordinate spaces. Then,
${\displaystyle\mathcal{L}_{\mathrm{clip}}({{\bm{A}}}_{V},{{\bm{A}}}_{T})=-\underset{({{\bm{x}}}_{V},{{\bm{x}}}_{T})\sim F_{\mathrm{clip}}}{\mathbb{E}}\bigg{[}({{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})^{T}{{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V}\bigg{]}}$
${\displaystyle+\underset{{{\bm{x}}}_{T}\sim F_{T}}{\mathbb{E}}\bigg{[}\log\bigg{(}\underset{{{\bm{x}}}_{V}\sim F_{V}}{\mathbb{E}}\big{[}\exp\big{(}({{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})^{T}{{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V}\big{)}\big{]}\bigg{)}\bigg{]}}$ (5)
Each coordinate space serves as an "augmentation" for the other coordinate space. Similar to InfoNCE, CLIP InfoNCE attract representations of ${{\bm{x}}}_{V}$ and ${{\bm{x}}}_{T}$ through first term, while regularizing with the Log-Sum-Exp term.
@@ 6.2 Results
Our main result for CLIPGMM states that we can learn a subset of Fisher subspace for the constituent modalities by minimizing the CLIP InfoNCE loss. Concretely,
@@
Suppose $\{w_{k},{\bm{\mu}}_{V,k},{\bm{\mu}}_{T,k},{\bm{\Sigma}}_{V},{\bm{\Sigma}}_{T}\}_{k\in[K]}$ is a CLIPGMM (Def 6.1). Let the Fisher subspace of $F_{V}$ be $S_{V,F}$ and $F_{T}$ be $S_{T,F}$. Denote ${{\bm{A}}}^{*}_{V},{{\bm{A}}}^{*}_{T}$ as the optimal solution of the CLIP InfoNCE loss (Eqn 6.1), i.e.,
${{\bm{A}}}^{*}_{V},{{\bm{A}}}^{*}_{T}=\underset{\begin{subarray}{c}{{\bm{A}}}_{V}\in\mathbb{R}^{d_{1}\times r}\\ {{\bm{A}}}_{T}\in\mathbb{R}^{d_{2}\times r}\end{subarray}}{\mathrm{argmin}}\ \mathcal{L}_{\mathrm{clip}}({{\bm{A}}}_{V},{{\bm{A}}}_{T})$
@@
The proof of Thm 6.2 can be found in Suppl E. In simple words, we learn the subset of the Fisher subspace instead of the exact space and hence is weaker result than the single-modal setting with augmentations. The result still shows that self-supervised learning for multi-modal data is better than spectral methods, since we are always a subset of $S_{F}$, which doesn’t hold for spectral methods (Fig 1).
The reason it is weaker than the single-modal GMM, however, is due to the fact that the means in the two spaces can vary arbitrarily. Hence, one can come up with adversarial set of means and covariances such that particular directions in the Fisher subspace (of either $F_{V}$ or $F_{T}$) doesn’t contribute to the inter-component distance and hence is not learned by the CLIP InfoNCE objective. For certain special configuration of model parameters such as ${\bm{\Sigma}}_{T}^{-\frac{1}{2}}{\bm{\mu}}_{T,k}={\bm{\Sigma}}_{V}^{-\frac{1}{2}}{\bm{\mu}}_{V,k}$, we can achieve $S_{T,\mathrm{clip}}=S_{T,F}$ and $S_{V,\mathrm{clip}}=S_{V,F}$.
@@
Figure 2: We empirically validate our theoretical findings for four main settings (see Sec 7). Fig (a), shows that self-supervised learning is robust to noise in Augmentation-enabled Distribution . Fig (b), shows that InfoNCE (and SimSiam) are invariant to variance orthogonal to fisher subspace. Fig (c), InfoNCE loss is better than spectral methods for every projection dimension, and Fig (d) shows that InfoNCE loss learns a good scaling within the fisher subspace
7 Synthetic Experiments
We validate our theoretical findings with experiments on synthetic data. We study the effect of noise $\delta$ in Augmentation-enabled Distribution and also how the rank $r$ of the projection matrix and condition number of covariance matrix effect the learned representations. Finally we show that apart from learning the optimal projection subspace, self-supervised learning also leads to a "good" scaling within the subspace.
@@ 7.1 Setup
We aim to evaluate different representation learning (linear dimensionality reduction) methods for the class of shared covariance GMMs (Eqn 3.1). We adopt a two-step process for evaluating the performance of the methods. The first step is dimensionality reduction with the target method, and the second step is clustering with an out-of-the-box clustering algorithm (i.e. K-Means  [^1] https://scikit-learn.org/1.5/modules/generated/sklearn.cluster.KMeans.html). We compare the methods based on their clustering performance in the second step.
Metrics
@@
Following prior work Jiang et al. (2020) we use Adjusted Rank Index (ARI) and Adjusted Mutual Information (AMI) to measure the quality of a clustering algorithms. We give the ground truth number of clusters to K-Means as input. ARI and AMI both vary between 0 and 1, with 0 corresponding to random clustering and 1 corresponding to perfect clustering.
Data Generation
@@
For all our experiments we consider $K=10$ equally likely (i.e. $w_{i}=1/K$) component mixtures. See Suppl G.1 for details.
Methods
@@
We consider four main baselines. 1) Ambient is clustering in ambient dimension, i.e. no mapping. 2) Random projects onto a random subspace of dimension $r$. 3) Optimal projects onto top $r$ directions in Fisher subspace. 4) PCA projects onto the top $r$ principal component of the samples (equivalent to SVD-subspace projection). On top of the baselines we also consider learning the mapping matrix ${{\bm{A}}}$ using gradient descent with respect to 5) InfoNCE loss and 6) SimSiam loss. Note that InfoNCE and SimSiam also learn scaling within the subspace (while baselines like PCA only project).
7.2 Results
@@
We observe that InfoNCE often performs better than Optimal (i.e. projection onto the Fisher subspace). We interpret that InfoNCE loss learns to scale within the subspace leading to better clustering performance (we present plots on ortho-normalized InfoNCE/SimSiam in Suppl G.2). Also observe SimSiam shows a sub-par performance compared to InfoNCE contrary to our theoretical results. We believe this is due the difficulty in optimizing the SimSiam objective (projected gradient descent) compared to InfoNCE.
Effect of noise in augmentations
@@
We vary the noise parameter $\delta$ for the AeD with everything else held constant. We can see in Fig 4(a) that InfoNCE is robust to the variation in $\delta$ and performs well even for small values of $\delta$.
Effect of variance orthogonal to $S_{F}$
@@
Here we consider the effect of increasing the variance in directions orthogonal to the Fisher subspace (i.e. making the pancake in Fig 1 flatter). We quantify the “flatness” as $\sigma_{\mathrm{min}}({\bm{\Sigma}})/\sigma_{\mathrm{max}}({\bm{\Sigma}})$. Fig 4(b) shows that InfoNCE is (almost) invariant to “flatness”, while PCA and Ambient fail when the variance orthogonal to Fisher subspace is large (i.e. $\sigma_{\mathrm{min}}(\Sigma)/\sigma_{\mathrm{max}}(\Sigma)$ is small).
Effect of projection dimension
@@
We vary the dimension of the lower dimensional space that we aim to learn. InfoNCE’s performance increases with increasing $r$ and finally plateaus.
Effect of variance within $S_{F}$
@@
We also show that InfoNCE learns a "good" scaling within the Fisher subspace. For this, we let the dimension of ambient space equal to the dimension of mean subspace (i.e. $d=K-1$). We select $K/2=5$ orthogonal directions and increase the variance in the subspace spanned by these direction by a factor of $\sigma_{\mathrm{max}}(\Sigma)/\sigma_{\mathrm{min}}(\Sigma)$. Since none of the baselines learn scaling in the subspace (and $d=r=K-1$), they perform almost the same, while InfoNCE improves.
8 Conclusion and Discussion
We study self-supervised learning in a classic setting of dimensionality reduction for Gaussian Mixture Models. A key idea that enabled this was defining the notion of augmentations and point pairs. We showed that for the problem of dimensionality reduction, if given the luxury of augmentations, we can go meaningfully beyond what is possible with completely unsupervised methods. Our main result underline that the contrastive learning paradigm learns the Fisher-optimal subspace for the class of shared-covariance GMMs. For future work, we hope to make use of our novel setting to design new, provable algorithms for learning Fisher-optimal subspaces.
Acknowledgments
Ali Kavis is funded in part by the Swiss National Science Foundation (SNSF) under grant number P500PT_217942.

### v3

@@ 5 Multi-modal Gaussian mixture models
For instance, each sample could be a pair of an image with its corresponding caption (for CLIP) or a search query with a relevant document that answers the query (for DPR). The objective in this multi-modal setup is to learn a joint representation space for both modalities. These joint representations can be used downstream for finding similarities between data points belonging to different ambient spaces, and are surprisingly competitive with fully supervised representations [21]. Our goal is to theoretically analyze the representations learned by such model, specifically the CLIP model. Next, we define CLIP-GMM to capture multi-modal data in a theory-friendly setting.
@@
(CLIP-GMM) A CLIP Gaussian mixture (CLIP-GMM) is defined as the probability distribution $F_{\mathrm{clip}}=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{V,k},{\bm{\Sigma}}_{V})\times\mathcal{N}({\bm{\mu}}_{T,k},{\bm{\Sigma}}_{T})$ where $w_{k}$ are the mixture weights, $\{{\bm{\mu}}_{V,k}\}_{k\in[K]},{\bm{\Sigma}}_{V}$ and $\{{\bm{\mu}}_{T,k}\}_{k\in[K]},{\bm{\Sigma}}_{T}$ are the parameters for the two coordinate spaces.
CLIP-GMM is a mixture of product distribution over Gaussians. To elaborate, sampling a pair $({{\bm{x}}}_{V},{{\bm{x}}}_{T})\sim F_{\mathrm{clip}}$ is a two step process. We first sample an underlying component index, and then draw independent samples from the component with the same index in the respective coordinate space. $F_{\mathrm{clip}}$ can be used to define marginal distribution over each coordinate space, for instance, $F_{V}=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{V,k},{\bm{\Sigma}}_{V})$.
@@ 5.1 Optimization objective
Following CLIP [21], we want to learn different representation functions for each modality. We let these functions be linear mappings, i.e., ${{\bm{A}}}_{V}\in\mathbb{R}^{d_{1}\times r},{{\bm{A}}}_{T}\in\mathbb{R}^{d_{2}\times r}$. Recall that optimal mapping matrices for each coordinate space would have their column space equal to that of Fisher subspace for the marginal distributions $F_{V},F_{T}$. Concretely, let the representation of a sample $({{\bm{x}}}_{V},{{\bm{x}}}_{T})$ be given by $({{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V},{{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})$, where ${{\bm{A}}}_{V},{{\bm{A}}}_{T}$ are the mappings. Then, we define CLIP InfoNCE loss [21] as,
$\mathcal{L}_{\mathrm{clip}}=\underset{({{\bm{x}}}_{V},{{\bm{x}}}_{T})\sim F_{\mathrm{clip}}}{\mathbb{-E}}\left[({{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})^{T}{{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V}\right]+\underset{{{\bm{x}}}_{T}\sim F_{T}}{\mathbb{E}}\Big[\log\Big(\underset{{{\bm{x}}}_{V}\sim F_{V}}{\mathbb{E}}\left[\exp\left(({{\bm{A}}}^{T}_{T}{{\bm{x}}}_{T})^{T}{{\bm{A}}}^{T}_{V}{{\bm{x}}}_{V}\right)\right]\Big)\Big]$
Each coordinate space serves as augmentation for the other. Similar to InfoNCE, CLIP InfoNCE attracts embeddings of ${{\bm{x}}}_{V}$ and ${{\bm{x}}}_{T}$ via the first term, while regularizing with the Log-Sum-Exp term.
@@ 5.2 Results
Our main result for CLIP-GMM states that we can learn a subset of Fisher subspace for the constituent modalities by minimizing the CLIP InfoNCE loss. We state our results in the following theorem.
@@
Suppose $\{w_{k},{\bm{\mu}}_{V,k},{\bm{\mu}}_{T,k},{\bm{\Sigma}}_{V},{\bm{\Sigma}}_{T}\}_{k\in[K]}$ is a CLIP-GMM. Let the Fisher subspace of $F_{V}$ be $S_{V,F}$ and $F_{T}$ be $S_{T,F}$. Denote ${{\bm{A}}}^{*}_{V},{{\bm{A}}}^{*}_{T}$ as the optimal solution of the CLIP InfoNCE loss,
${{\bm{A}}}^{*}_{V},{{\bm{A}}}^{*}_{T}=\underset{\begin{subarray}{c}{{\bm{A}}}_{V}\in\mathbb{R}^{d_{1}\times r},\ {{\bm{A}}}_{T}\in\mathbb{R}^{d_{2}\times r}\end{subarray}}{\mathrm{argmin}}\ \mathcal{L}_{\mathrm{clip}}({{\bm{A}}}_{V},{{\bm{A}}}_{T})$
@@
The proof of Theorem˜5.2 can be found in Appendix˜E. In simple words, CLIP InfoNCE learns the subset of the Fisher subspace instead of the exact space, which is weaker than the single-modal setting. This is due to the fact that the means in the two spaces can vary arbitrarily. Hence, one can choose means and covariances adversarially such that particular directions in the Fisher subspace of either $F_{V}$ or $F_{T}$ does not contribute to the inter-component distance and are not learned by the CLIP InfoNCE. For certain special configuration of model parameters, such as ${\bm{\Sigma}}_{T}^{-\frac{1}{2}}{\bm{\mu}}_{T,k}={\bm{\Sigma}}_{V}^{-\frac{1}{2}}{\bm{\mu}}_{V,k}$, we can achieve $S_{T,\mathrm{clip}}=S_{T,F}$ and $S_{V,\mathrm{clip}}=S_{V,F}$. The result still shows that self-supervised learning for multi-modal data is better than non-supervised methods. Accessing augmentations are strictly weaker than the knowledge of the labels, and they also naturally occur in the multi-modal setting. The fact that contrastive losses learn a subset of the optimal subspace verifies their capabilities.
@@
Figure 2: We empirically validate our theoretical findings for four main settings. (a) InfoNCE is robust to noise in augmentations for various values of $\delta$. (b) InfoNCE (and SimSiam) are invariant to variance orthogonal to Fisher subspace. (c) InfoNCE outperforms spectral methods for every mapping dimension. (d) InfoNCE learns a good scaling within the Fisher subspace
6 Experiments
We validate our theoretical findings with experiments on synthetic and real data. For the synthetic data, we study the effect of noise $\delta$ in Augmentation-enabled Distribution , rank $r$ of the projection matrix and condition number of covariance matrix on learned representations. For the real data experiments, we compare self-supervised methods against baselines for clustering CIFAR100 dataset on low dimensional subspaces using K-Means.
@@ 6.1 Setup
We evaluate different linear dimensionality reduction methods for SharedGMMs. We adopt a two-step process; the first step is dimensionality reduction with the target method, and the second step is clustering with an out-of-the-box clustering algorithm (i.e. K-Means ). We compare the methods based on their clustering performance in the second step using well-known metrics.
@@
Metrics. Following prior work [14] we use Adjusted Rank Index (ARI) and Adjusted Mutual Information (AMI) to measure the quality of a clustering algorithms. We give the ground truth number of clusters to K-Means as input. ARI and AMI both vary between 0 and 1, with 0 corresponding to random clustering and 1 corresponding to perfect clustering.
@@
Data generation. For synthetic experiments we consider $K=10$ with equally likely component mixtures. For the CIFAR100 experiments, we randomly sample $K=20$ classes out of 100. See Suppl G.1 for details.
@@
Methods. We consider learning a mapping matrix $\bm{A}$}} using gradient descent with respect to InfoNCE and SimSiam loss and we compare them againts 5 baselines. Ambient is clustering in ambient dimension. Random projects onto a random $r$-dimensional subspace. Optimal projects onto top $r$ directions in Fisher subspace. PCA projects onto $r$-dimensional $S_{SVD}$. LDA finds a projection that maximizes inter-class variance with respect to intra-class variance (Optimal in the synthetic case).
6.2 Synthetic experiments
@@
We observe that InfoNCE often performs better than Optimal, projection onto the Fisher subspace. We interpret that InfoNCE loss learns to scale within the subspace leading to better clustering performance (see App G.2 for orthonormalized plots of InfoNCE/SimSiam). Moreover, SimSiam shows a sub-par performance compared to InfoNCE. We believe this is due the difficulty in optimizing the SimSiam objective (projected gradient descent) compared to InfoNCE.
@@
Effect of noise in augmentations. We vary the noise parameter $\delta$ for the AeD with everything else held constant. We can see in Fig 4(a) that InfoNCE is robust to the variation in $\delta$.
@@
Effect of variance orthogonal to $S_{F}$. Increasing the variance orthogonal to the Fisher subspace is equivalent to making the pancakes in Fig 1 flatter. We quantify “flatness” as $\sigma_{\mathrm{min}}({\bm{\Sigma}})/\sigma_{\mathrm{max}}({\bm{\Sigma}})$. Fig 4(b) shows that InfoNCE is (almost) invariant to “flatness”, while PCA and Ambient fail when the variance orthogonal to Fisher subspace is large (i.e. $\sigma_{\mathrm{min}}(\Sigma)/\sigma_{\mathrm{max}}(\Sigma)$ is small).
@@
Effect of projection dimension. We vary the dimension of the lower dimensional space that we aim to learn. InfoNCE’s performance increases with increasing $r$ and finally plateaus.
@@
Effect of variance within $S_{F}$. We show that InfoNCE learns a "good" scaling within the Fisher subspace. We set the dimension of ambient space equal to the dimension of mean subspace (i.e. d=K1d=K-1). We select $K/2$ orthogonal directions and increase the variance in the subspace spanned by these direction by a factor of $\sigma_{\mathrm{max}}(\Sigma)/\sigma_{\mathrm{min}}(\Sigma)$. Since none of the baselines learn scaling in the subspace, they perform almost the same, while InfoNCE outperforms all the baselines.
6.3 Real-data experiments on CIFAR-100
We consider images from CIFAR-100 as inputs instead of generating data from a synthetic GMM. We consider 20 fine-labeled classes (10K images) where each class corresponds to a single component. We convert images into a 256-dimensional vector by subsampling, grayscaling, mean scaling. Images belonging to the same class will be the augmentations of each other. We still consider a linear model and hence have the same baselines. We measure the clustering performance using ARI and AMI on mapping to $r=5,10,15,19$-dimensional subspaces in Table˜1. Surprisingly, InfoNCE could still outperform LDA in certain metrics even in the real data setting where naturally, our data distribution assumptions and augmentation will no longer hold. LDA has the best performance for all dimensions in terms of ARI (first value in each column) which indicates that it excels at maintaining local groupings. On the other hand, InfoNCE achieves larger AMI values underlining that contrastive learning could be better at preserving overall class distributions.
Table 1: Clustering performance of linear dimensionality reduction methods on CIFAR100: Bolded and underlined values indicate the best and second-best scores for each column. We report clustering performance for K-Means on $r=5,10,15,19$-dimensional subspaces using the linear mappings learned by 6 different methods. We measure the clustering performance using ARI and AMI, which is reported as pairs, in the format ARI | AMI. InfoNCE and LDA outperform the rest by significant margins in all mapping dimensions. LDA shows the best performance in terms of ARI, while InfoNCE achieves higher AMI values across the board.
Mapping Dimension
Method 5 dim 10 dim 15 dim 19 dim 30 dim
Random 0.01475 | 0.04694 0.01567 | 0.05241 0.01894 | 0.05796 0.02446 | 0.07188 0.02503 | 0.07856
PCA 0.03360 | 0.09254 0.03292 | 0.09509 0.03435 | 0.09427 0.03154 | 0.09269 0.03361 | 0.09562
Ambient 0.03160 | 0.09481 0.03160 | 0.09481 0.03160 | 0.09481 0.03160 | 0.09481 0.03160 | 0.09481
SimSiam 0.02870 | 0.09581 0.02732 | 0.09114 0.03707 | 0.10830 0.02868 | 0.09498 0.03562 | 0.10540
InfoNCE 0.05065 | 0.12730 0.05138 | 0.12801 0.05402 | 0.12835 0.05285 | 0.12947 0.05182 | 0.12548
LDA 0.05067 | 0.12360 0.05489 | 0.11782 0.06352 | 0.12359 0.05970 | 0.11933 0.05970 | 0.11933
6.4 Real-data experiments on ImageNet
We consider images from ImageNet as inputs instead of generating data from a synthetic GMM. We consider 20 fine-labeled classes (10K images) where each class corresponds to a single component, and images belonging to the same class serve as the augmentations. Since image classification is highly non-linear, we construct a setup to map them to a linearly seperable space. For this we represent each image using ResNet-50’s last layer representations (i..e, 2048 dimensional vectors). We measure the clustering performance using ARI and AMI on mapping to $r=5,10,15,19$-dimensional subspaces in Table˜2. We can see that InfoNCE consistently outperforms LDA in this setting.
Table 2: Clustering performance of linear dimensionality reduction methods on ImageNet: Bolded and underlined values indicate the best and second-best scores for each column. We report clustering performance for K-Means on $r=5,10,15,19$-dimensional subspaces using the linear mappings learned by 6 different methods. We measure the clustering performance using ARI and AMI, which is reported as pairs, in the format ARI | AMI. InfoNCE and LDA outperform the rest by significant margins in all mapping dimensions. LDA shows the best performance in terms of ARI, while InfoNCE achieves higher AMI values across the board.
Mapping Dimension
Method 5 dim 10 dim 15 dim 19 dim 30 dim
Random 0.03026 | 0.08296 0.07231 | 0.14005 0.13140 | 0.23209 0.14405 | 0.23363 0.26177 | 0.38685
PCA 0.23408 | 0.46447 0.39843 | 0.59117 0.49903 | 0.65898 0.50954 | 0.66159 0.56355 | 0.70163
Ambient 0.48641 | 0.67233 0.48641 | 0.67233 0.48641 | 0.67233 0.48641 | 0.67233 0.48641 | 0.67234
SimSiam 0.37581 | 0.58233 0.56084 | 0.70246 0.58705 | 0.72227 0.60259 | 0.73159 0.62970 | 0.74476
InfoNCE 0.84451 | 0.88831 0.98182 | 0.98115 0.92963 | 0.97390 0.99621 | 0.99579 0.93317 | 0.97721
LDA 0.47112 | 0.69828 0.66967 | 0.83895 0.82550 | 0.89103 0.94744 | 0.94990 0.94745 | 0.94990
7 Conclusion and limitations
We study contrastive learning in the classical setting of linear dimensionality reduction for GMMs for which we define a generalized notion of imperfect augmentations. Our main result underlines that the contrastive methods learn the Fisher-optimal subspace for the class of shared-covariance GMMs; going beyond the capabilities of unsupervised methods and matching that of fully supervised strategies. Our work particularly focuses on linear mappings, and it is an important open problem to theoretically verify performance of contrastive methods when projectors are non-linear, which is usually the case in practice. Similarly, we acknowledge that our results are not immediately generalizable for data distributions beyond GMMs, which we will investigate as future work.